OpenStudy (maddytarin_98):

The solutions to the equation2x^2+-1=-3/2 or x=____

OpenStudy (maddytarin_98):

really need help, taking timed quiz and looking for answers! much appreciated

MARC (marc_d):

First,combine the like terms

OpenStudy (rolldextro):

\[\frac{ i }{ 2 } an d \frac{ -i }{ 2 }\]

OpenStudy (rolldextro):

Solve for x: 2 x^2-1 = -3/2 Add 1 to both sides: 2 x^2 = -1/2 Divide both sides by 2: x^2 = -1/4 Take the square toor

OpenStudy (rolldextro):

\[\sqrt{x}=\sqrt{\frac{ -1 }{ 4 }}\]